OpenStudy (anonymous):

find the equation of the line (5,1)

Directrix (directrix):

Infinitely many lines can pass through a single point. Is there more information about the problem?

Directrix (directrix):

Thanks for the diagram post. The line passes through the origin (0,0) So, the task is to find the equation determined by the points (0,0) and (5,1). So....

Directrix (directrix):

y = (1/5) x http://www.webmath.com/equline1.html

Directrix (directrix):

Questions?

OpenStudy (anonymous):

so the answer is 1/5

OpenStudy (anonymous):

The ordered pair (2, 36) is a solution for which equation(s)? Check all that apply.

OpenStudy (anonymous):

A. y = (3x)2 B. y = 2x2 C. y = 7x2 + 2 D. y = (x + 4)2